OpenStudy (anonymous):

Fundamental Theorem of Calculus help needed: Find the numbers "b" such that the average value of f(x)=2 + 7x - x^3 on the interval [0,b] = 3

OpenStudy (phi):

The idea is that a "box" from x=0 to b, with a height of 3 has the same area as the area under the curve from x=0 to b

OpenStudy (phi):

The area under the curve is found by integrating \[ \int_0^b 2 + 7x - x^3 \] and you want this area to be the same area as the box with width (b-0)*3 or 3b \[ 3b= \int_0^b 2 + 7x - x^3 \] The first step is integrate the right hand side. then solve for b

OpenStudy (anonymous):

So the integration would be 2x + 7x^2/2 - x^4/4

OpenStudy (phi):

yes, now evaluate between the bounds 0 to b you get 2b + (7/2)b^2 - (¼) b^4 = 3b

OpenStudy (anonymous):

Okay Okay Then aI can subtract 2b to get 3.5b^2 - b^4/4 = 1b Where am I left then?

OpenStudy (phi):

you get b=0 and 0.25 b^3 -3.5b + 1 =0 this is difficult to solve by hand, but wolfram gives the values.

OpenStudy (anonymous):

Oh of course you do. Thank you I will look there.